A rectangle has 200 feet of fencing with which to enclose two adjacent rectangle corrals. What dimensions should be used so the enclosed area will be a maximum? what is the maximum area?
Answer by stanbon(75887)

A rectangle has 200 feet of fencing with which to enclose two adjacent rectangle corrals. What dimensions should be used so the enclosed area will be a maximum? what is the maximum area?
-------------
Draw the picture of 2 adjacent rectangles.
Let height be "h" (Note: there are 3 of them)
Then width = (1/2)(200-3h)
---------------
Area = (h/2)(200-3h)
Area = 100h - (3/2)h^2
-----
Max area occurs when
h = -b/(2a)
= -100/(2(-3/2))
= 100/3 ft.
--------------------------
Then length = (1/2)(200-(3/2)(100/3))
= (1/2)(50)
= 25 ft.
==================
Max area = 25*(100/3) = 833 1/3 sq. ft.
===========================================
Cheers,
Stan H.